OpenStudy (phagey):

A store manager uses a markup rate of 23% on all apparel. Find the cost of a shirt that sells for $12.99. Use the formula S = C + R x C, where S is the selling price, C is the cost, and R is the markup rate. $15.98 $2.99 $10.56 $56.48

OpenStudy (phagey):

@welshfella

zepdrix (zepdrix):

\[\large\rm S=C+R\cdot C\] They've provided us with the sale price \(\large\rm S=12.99\) and the markup rate \(\large\rm C=.23\) (This is 23% converted to a decimal value).

zepdrix (zepdrix):

We'll plug this information in,\[\large\rm 12.99=C+.23C\]and attempt to solve for C, the cost of the shirt.

zepdrix (zepdrix):

You have `like-terms` on the right side of the equation. So how much is 1C added to .23C ?

OpenStudy (phagey):

ok

OpenStudy (phagey):

is it 10.56?

zepdrix (zepdrix):

12.99 = 1.23C So then C = 12.99/1.23 = 10.56 Yay good job \c:/